OpenStudy (anonymous):

5x^2 - 6x = 1 using complete sentences explain who used a better method completing the square or factoring would be better?

OpenStudy (anonymous):

I'm sorry guys I need answers right now

OpenStudy (lgbasallote):

factoring would be easier but completing the square is safer

OpenStudy (aravindg):

yep i agree with @lgbasallote

OpenStudy (king):

But then if u were in an exam factoring would be fast and easy...

OpenStudy (aravindg):

yep always try to think !! i am sure u will then see which is the factors easily

Parth (parthkohli):

I think completing the square, because I'm no able to find two numbers that multiply to get -5 and add to get -6. :/

OpenStudy (anonymous):

thanks!

OpenStudy (lgbasallote):

how about -5 and -1 @ParthKohli

OpenStudy (anonymous):

Excuse me, but this doesn't factor well. What context are we talking here?

Parth (parthkohli):

-5 * -1 = 5

OpenStudy (lgbasallote):

what do you guys mean doesnt factor out well?? 5x^2 - 6x - 1= 0 5x^2 - 5x - x -1 = 0 it's factorable o.O

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5x(x - 1) - (x + 1) = 0 }\) NOPE IT ISNT

OpenStudy (anonymous):

I mean that it does not factor over the integers. This is the usual subset of R chosen.

OpenStudy (lgbasallote):

hmm you have a point

Parth (parthkohli):

As I said, \(-5 \times -1= 5\)

OpenStudy (anonymous):

Katie was given the incorrect answer.

Parth (parthkohli):

And lgba was given an undeserved medal.

OpenStudy (anonymous):

Yes, but that's not his fault. Everyone is wrong from time to time. :P